# Explaining the Implicit Neural Canvas: Connecting Pixels to Neurons by Tracing their Contributions

### v1

@@ 4.3 Contribution is not Intensity
While there is naturally some intensity bias, as a result of the pixelwise loss functions used by these networks, this does not fully explain the behavior of the representation. To discover other concepts that the INR potentially pays attention to, we subtract the intensity image from the contribution heat map. In Figure 5, this reveals that in addition to intensity, the INR layers also pay attention to other low-level information such as edges and textures. We note that the NeRV layers and the early FFN layers have a tendency to over-represent the people in the second image, relative to intensity, while the FFN layer 3 contributions are largely consistent with intensity. Additionally, not all edges are treated equally, with NeRV having large contributions for some edges, and small contributions for others.
@@ 4.4 Representation is Distributed
When training an INR, one might naturally wonder, since there are more pixels than neurons, how many pixels are represented by each neuron, and if this representation tends to be distributed. To analyze the manner in which neuron-to-pixel contributions are distributed, we compute the ratio of a neuron’s contribution to a pixel to the total contribution across neurons at a pixel. We threshold on this ratio to determine whether a pixel receives a significant portion of its contribution from a specific neuron or not. The threshold used is a simple reciprocal of the number of neurons in a layer. We then compute the total number of pixels that a specific neuron contributes to meaningfully and sort this across neurons, repeating this analysis individually for different layers of the FFN and NeRV.
@@
We show these contribution counts in Figure 6. Since different layers have different numbers of neurons, we resample the neurons of each layer to a common length using linear interpolation to preserve their trends. Notice how the outermost layers of the FFN and NeRV (layer 3 and head layer respectively) show the highest trends with a decreasing trend towards the inner layers of both networks. Note that nearly $50\%$ of the neurons in FFN Layer 2 have almost no contributions, producing a flat line in the first half. This suggests that INRs follow similar trends to discriminative networks, where earlier layers contribute towards structural features while later layers have more stochastic features.
@@
To continue our exploration of how the learned INRs distribute the image representations, we compute at each spatial image location, the density of neurons that contribute to it above a threshold, $\tau$. To set $\tau$ for a certain layer, consider the sorted raw contributions of all neurons across locations of the image. We find that a large portion of the contributions are small and the number of larger contributions is fewer in comparison (see Appendix). We obtain the raw contribution value that lies at the $x$-th percentile of this curve, i.e., the contribution of the neuron to the left of which lies $x\%$ of the area under the curve. The value corresponding to this percentile is chosen as the threshold for all neuron contribution maps. We show this for two different thresholds in Figure 7, which shows the regions of the image with a higher density of dedicated neurons.
@@
High intensity and highly textured areas not only have high contributions as seen earlier in Figure 5, but they also have a large portion of neurons dedicated to them. Increasing the threshold reveals the image areas with the highest neuron density, and this reveals an overwhelming bias for edges with NeRV. However, all these trends are greatly subdued for the FFN at both its head and penultimate layers, suggesting that while the magnitude of its contributions correlates almost perfectly with the image intensity, its actual representation is more evenly distributed by comparison.
@@ 4.5 Objects, Categories, and Motion
To analyze whether INRs have meaningful representations in terms of objects, we aggregate neuron contributions over pixels in each instance’s segmentation map. Consider the two sample videos shown in Figure 8. Each has a different set of objects with some categories having multiple instances. Instances may move, disappear and reappear over time. For the FFN, we sample three neurons each from the three layers and depict each neuron’s normalized contributions to every instance (as a percentage of contributions to all instances) in the first frame of the video. Notice how contributions to instances of the same category are not grouped together. This is related our findings in Figure 4, and motivates our claim that these networks learn low-level object features, but lack class-level semantics.
For the NeRV, we sample two neurons each from the head layer Block 3, and track the normalized per-instance contributions over multiple frames of the video. We see that a neuron’s contributions to an instance remains relatively static over time. When there is fluctuation, we notice that this is often correlated with instances that move, such as persons. Another thing to note is how contributions to high intensity instances such as sky are higher, which is in agreement with our earlier observations.
@@
Further delving into how neurons in NeRV respond to motion, we visualize the optical flow map and the fluctuation of total contribution in the last 2 layers of a NeRV. We are careful to select videos with no camera motion, which could lead to spurious or whole-frame flow. Figure 9 depicts this for two different time steps in each video (the flow and contribution fluctuation are computed between a specific frame and the immediately preceding frame). We see that the changes in contributions are focused more in areas around the regions of flow. It appears that when a moving object reveals new background, this causes fluctuations in the spatially proximal pixels, even though they technically contain no motion themselves. Thus, we see that in spite of its lack of high-level object semantics, the changes in how a NeRV represents a video across time are dominated by the motion of the entities in the video.
@@ 4.6 Neurons can be Clustered by Contributions
We propose a unique approach for computing a vector representation of INR neurons. Inspired by the relatively high variance for Gabor clusters in Figure 4, we take the pixels in each frame of each video, and cluster them according to Gabor filter features with $k$ clusters. We then aggregate the contribution map values according to the clusters, yielding a $k$ dimensional vector representation of every neuron. We can then project these representations with UMAP to allow for plotting them in 2D space, which we do in Figure 10. As this Figure shows, there are many different “types” of neurons when we separate them according to edge contributions. This representation mechanism we develop with XINC gives another powerful method for visualizing and examining the internals of INR.
@@ 5 Conclusion
In this paper, we propose XINC, a novel framework for analyzing INRs for image and video representation. We use XINC to find interesting trends with CNNs and MLPs, such as how while they lack high-level object semantics, they do have biases that mimic object representation, particularly with edges and colors. We also show how fluctuations in contribution for NeRV are driven by motion. We show how XINC can help us identify interesting groups of related neurons. We provide XINC, which is flexible for MLPs and CNNs, as a resource to help drive future research for INRs.
@@
The main purpose of this appendix is to expand our results from the main paper by showing analysis for frames from other videos, to show that our analysis and findings hold across a variety of different videos. For additional implementation details, including how we construct the contribution maps for various layers, please refer to our code.
6 Gabor features
Gabor filters are used to analyze different textures by detecting patterns at different orientations and scales. To extract Gabor features in our experiments, we employ Gabor filters at four distinct orientations and three different scales to account for a diverse range of textures. For further details on how we aggregate Gabor features for grouping contributions into Gabor clusters, please refer to our code.
7 Contribution Maps
@@
Using Gabor filter features, we cluster neurons at each layer of MLP-based and CNN-based INRs and plot them using UMAP. For a few of these clusters, Figures 11, 12, 13, 14 and 15 show the contribution maps of four neurons sampled from each cluster. We see that early FFN layers learn Fourier patterns, and the last layer of NeRV tends to resemble the image. Further, note how neurons belonging to the same cluster tend to learn similar contribution maps as compared to neurons belonging to different clusters.
@@
Figure 11: Neuron Clusters. Contribution maps of sample clustered neurons for the first (red-bordered) frame for video 1.
@@
Figure 12: Neuron Clusters. Contribution maps of sample clustered neurons for the first (red-bordered) frame for video 2.
Figure 13: Neuron Clusters. Contribution maps of sample clustered neurons for the first (red-bordered) frame for video 3.
Figure 14: Neuron Clusters. Contribution maps of sample clustered neurons for the first (red-bordered) frame for video 4.
Figure 15: Neuron Clusters. Contribution maps of sample clustered neurons for the first (red-bordered) frame for video 5.
@@
In Figure 16, we provide a supplement to Figure 10 by clustering neurons from models with five different seeds, for each layer of both MLP-based and CNN-based INRs on frames from five videos.
8 Grouping Contributions
In Figure 17, we provide a supplement to Figure 4 by showing results for frames from 5 additional videos. As expected, the major trends hold. We see far higher neuron contribution difference variances when using the instances, RGB clusters, and Gabor clusters, compared to space clusters (gridcells). We see that in general Gabor and RGB clusters have equal explanatory power for neuron contributions compared to instances, reinforcing our hypothesis that these networks have low-level, rather than high-level, object semantics. Interestingly, the trends are consistent across all layers, except for the trends for space are weaker with the NeRV than the FFN, perhaps due to the spatial bias from NeRV’s convolutional kernels.
@@ 9 Representation is Distributed
Figure 20: Distribution of Neuron Contributions. We plot the distribution over all neuron contributions in each layer. This shows that a large fraction of neuron contributions tend to be relatively smaller in magnitude. The contribution values corresponding to the $10^{th}$ and $50^{th}$ percentile of each distribution are used in selecting thresholds for Figures 7 and 19.
@@
Figure 20 shows how a large portion of the raw contributions are relatively smaller in magnitude for each layer. The thresholds for Figure 19 are selected using the contribution value at the $10^{th}$ and $50^{th}$ percentile of these curves.

### v2

@@ 4.3 Contribution is not Intensity
While there is naturally some intensity bias, as a result of the pixelwise loss functions used by these networks, this does not fully explain the behavior of the representation. To discover other concepts that the INR potentially pays attention to, we subtract the intensity image from the contribution heat map. In Figure 5, this reveals that in addition to intensity, the INR layers also pay attention to other low-level information such as edges and textures. We note that the NeRV layers and the FFN layer 2 have a tendency to over-represent the people in the second image, relative to intensity, while the FFN layer 3 contributions are largely consistent with intensity. Additionally, not all edges are treated equally, with NeRV having large contributions for some edges, and small contributions for others.
@@ 4.4 Representation is Distributed
When training an INR, one might wonder, since there are more pixels than neurons, how many pixels are represented by each neuron, and if this representation tends to be distributed. To analyze the manner in which neuron-to-pixel contributions are distributed, we compute the ratio of a neuron’s contribution to a pixel to the total contribution across neurons at a pixel. We threshold on this ratio to determine whether a pixel receives a significant portion of its contribution from a specific neuron or not. The threshold used is a simple reciprocal of the number of neurons in a layer. We then compute the total number of pixels that a specific neuron contributes to meaningfully and sort this across neurons, repeating this analysis individually for different layers of the FFN and NeRV.
@@
We show these contribution counts in Figure 6. Since different layers have different numbers of neurons, we resample the neurons of each layer to a common length using linear interpolation to preserve their trends. Notice how the outermost layers of the FFN and NeRV (layer 3 and head layer respectively) show the highest trends with a decreasing trend towards the earlier layers of both networks. Note that nearly $50\%$ of the neurons in FFN Layer 2 have almost no contributions, producing a flat line in the first half. The presence of these dead neurons in this and other layers points to the suitability of these networks for data compression, and helps explain why extensive model pruning can be done with limited impact on reconstruction quality [7].
@@
To further understand how the learned INRs distribute the image representations, we compute at each spatial image location, the density of neurons that contribute to it above a threshold, $\tau$. To set $\tau$ for a certain layer, consider the sorted raw contributions of all neurons across locations of the image. We find that a large portion of the contributions are small and the number of larger contributions is fewer in comparison. For an illustration of this, see Figure 18. We obtain the raw contribution value that lies at the $x$-th percentile of this curve, i.e., the contribution of the neuron to the left of which lies $x\%$ of the area under the curve. The value corresponding to this percentile is chosen as the threshold for all neuron contribution maps. We show this for the $10^{\text{th}}$ and $50^{\text{th}}$ percentile thresholds in Figure 7, which shows the regions of the image with a higher density of dedicated neurons.
@@
High intensity and highly textured areas not only have high contributions as seen earlier in Figure 5, but they also have a large portion of neurons dedicated to them. Increasing the threshold highlights the image areas with the highest neuron density, and this reveals an overwhelming bias for edges with NeRV. However, all these trends are greatly subdued for the FFN at both its head and penultimate layers, suggesting that while the magnitude of its contributions correlates almost perfectly with the image intensity, its actual representation is more evenly distributed by comparison.
@@ 4.5 Objects, Categories, and Motion
To analyze whether INRs have meaningful representations in terms of objects, we aggregate neuron contributions over pixels in each instance’s segmentation map. Consider the two sample videos shown in Figure 8. Each has a different set of objects with some categories having multiple instances. Instances may move, disappear and reappear over time. For the FFN, we sample three neurons each from the three layers and depict each neuron’s normalized contributions to every instance (as a percentage of contributions to all instances) in the first frame of the video. Notice how contributions to instances of the same category are not grouped together. This is related to our findings in Figure 4, and motivates our claim that these networks learn low-level object features, but lack class-level semantics.
For NeRV, we sample two neurons each from the head layer Block 3, and track the normalized per-instance contributions over multiple frames of the video. We see that a neuron’s contributions to an instance remains relatively static over time. When there is fluctuation, we notice that this is often correlated with instances that move, such as persons. Another thing to note is how contributions to high intensity instances such as sky are higher, which is in agreement with our earlier observations.
@@
Further delving into how neurons in NeRV respond to motion, we visualize the optical flow map and the fluctuation of total contribution in the last 2 layers of a NeRV. Figure 9 depicts this for two different time steps in each video (the flow and contribution fluctuation are computed between a specific frame and the immediately preceding frame). We see that the changes in contributions are focused more in areas around the regions of flow. It appears that when a moving object reveals new background, this causes fluctuations in the spatially proximal pixels, even though they technically contain no motion themselves. Thus, we see that in spite of its lack of high-level object semantics, the changes in how a NeRV represents a video across time are dominated by the motion of the entities in the video.
@@ 4.6 Neurons can be Clustered by Contributions
We propose a unique approach for computing a vector representation of INR neurons. Inspired by the relatively high variance for Gabor clusters in Figure 4, we take the pixels in each frame of each video, and cluster them according to Gabor filter features with $k$ clusters. We then aggregate the contribution map values according to the clusters, yielding a $k$ dimensional vector representation of every neuron. We can then project these representations with UMAP to allow for plotting them in 2D space, which we do in Figure 10. As this figure shows, there are many different “types” of neurons when we separate them according to edge contributions. This representation mechanism we develop with XINC provides another powerful method for visualizing and examining the internals of INR.
@@ 5 Conclusion
In this paper, we propose XINC, a novel framework for analyzing INRs for image and video representation. Using XINC, we find interesting trends with CNNs and MLPs, such as how while they lack high-level object semantics, they do have biases that mimic object representation, particularly with edges and colors. We also show how fluctuations in contribution for NeRV are driven by motion. We show how our framework can help us identify interesting groups of related neurons. We provide both the framework and these findings to help drive future research for INRs.
Acknowledgements. We would like to thank Soumik Mukhopadhyay for his helpful feedback while we prepared the manuscript. This project was partially funded by NSF CAREER Award (#2238769) to AS.
@@
The main purpose of this appendix is to expand our results from the main paper by showing analysis for frames from additional videos, to demonstrate that our analysis and findings hold across a variety of different videos.
6 Implementation Details
6.1 INR Architecture Settings
Neither the INR or NeRV have established out-of-the-box settings for arbitrary, in-the-wild videos. Thus, we have to carefully explore and select settings that allow for meaningful comparisons. We try to ensure the FFN and NeRV have similar compression ratios; however, this is not fully practical for one key reason. NeRV is able to leverage the temporal redundancies in addition to the spatial redundancies. We also try to ensure that both networks achieve similar reconstruction quality. For the videos we choose, the NeRVs have a mean PSNR of 33.50, and for the frames, the FFNs have a mean PSNR of 34.29.
For all the NeRVs, we ensure they have 978,557 parameters; for all FFNs, we ensure they have 32,971; for a 30 frame video (for Cityscapes-VPS all are 30 frames, for VIPSeg we only use videos between 30 and 45 frames), where the NeRV represents all 30 frames, and a given FFN represents a single frame, this gives roughly equivalent bits-per-pixel. For the FFN, we feed the 208-dimension Fourier position encoding to a network with 3 layers, with hidden sizes of 104 for the inner layers, ReLU activations, and output as 3-channel rgb predictions. For the NeRV, we use 4 layers, with the upsampling layers having strides 4, 2, and 2, respectively. We borrow the parameter reduction trick from HNeRV to balance parameters between layers, and set this to $r=1.2$ [9], with a minimum width of 6, and the model size hyperparameter set to 1 million. The inner layers use GELU activations [20]. For other settings, we use the original NeRV defaults [7]. Both are trained for 1000 epochs with L2 loss.
6.2 Gabor Features
We already have mechanisms for addressing some types of low-level and high-level features. Since we have dense instance and background segmentation masks, we are able to identify how neuron contributions correspond with instances and background. We can also cluster pixels by color and space. However, we notice low-level patterns in the contribution maps, such as a tendency to focus on edges, that are not captured fully by instance masks, color, or space. So, we use an additional mechanism for clustering pixel locations – Gabor filters [16, 33].
Gabor filters offer a robust method for texture analysis, detecting patterns across various orientations and scales. In our experiments, we utilize Gabor filters with four orientations and three scales, ensuring a comprehensive analysis of diverse textures within the contribution maps. The filters are applied to the maps, producing a distinct feature map for each filter. These feature maps are then stacked. For each pixel coordinate, a feature vector is constructed from values across all filters. These vectors are utilized to group contributions with similar traits into Gabor clusters (see Section 4.6). Please refer to the code for precise settings and implementation.
7 Further Results
7.1 Contribution Maps
We show the missing layers from Figure 3 in Figure 11. Note that the earliest layers have patterns that heavily correlate with the fourier features (FFN) and positional encoding (NeRV). Also, due to the nature of the PixelShuffle, for the first layer of NeRV, some neurons cannot possibly represent certain pixels. In that sense, the representation of each neuron is forced to be somewhat sparse.
@@ 7.1 Contribution Maps
Using Gabor filter features, we cluster neurons at each layer of MLP-based and CNN-based INRs and plot them using UMAP. For a few of these clusters, Figures 12 and 13 show the contribution maps of four neurons sampled from each cluster. We see that early FFN layers learn Fourier patterns, and the last layer of NeRV tends to resemble the image. Further, note how neurons belonging to the same cluster tend to learn similar contribution maps as compared to neurons belonging to different clusters.
Figure 11: The implicit neural canvas. We show the contribution maps for sample first layer neurons of FFN and NeRV.
@@
Figure 12: Neuron Clusters. We cluster neurons for each layer/block into 4 clusters, and then sample 4 neurons from each cluster. We show contribution maps of sample clustered neurons for the first (red-bordered) frame for the indicated video (top). See another video in Figure 13.
@@
Figure 13: Neuron Clusters. Following the same procedure as in Figure 12, we cluster neurons for each layer/block into 4 clusters, and then sample 4 neurons from each cluster. We show contribution maps of sample clustered neurons for the first (red-bordered) frame for the indicated video (top).
@@
In Figure 14, we provide a supplement to Figure 10 by clustering neurons from models with five different seeds, for layers of both MLP-based and CNN-based INRs on frames from five videos.
7.2 Grouping Contributions
In Figure 16, we provide a supplement to Figure 4 by showing results for frames from 5 additional videos. As expected, the major trends hold. We see far higher neuron contribution difference variances when using the instances, RGB clusters, and Gabor clusters, compared to space clusters (gridcells). We see that in general Gabor and RGB clusters have equal explanatory power for neuron contributions compared to instances, reinforcing our hypothesis that these networks have low-level, rather than high-level, object semantics. Interestingly, the trends are consistent across all layers, except for the trends for space are weaker with the middle layers of NeRV than with the FFN, perhaps due to the spatial bias from NeRV’s convolutional kernels.
@@ 7.3 Representation is Distributed
Figure 18: Distribution of Neuron Contributions. We plot the distribution over all neuron contributions in each layer. This shows that a large fraction of neuron contributions tend to be relatively smaller in magnitude. The contribution values corresponding to the $10^{\text{th}}$ and $50^{\text{th}}$ percentile of each distribution are used in selecting thresholds for Figures 7 and 17.
Figure 19: Neuron contributions to things and stuff for the first two blocks of NeRV. See Figure 8 for other blocks/layers.
@@
Figure 18 shows how a large portion of the raw contributions are relatively smaller in magnitude for each layer. The thresholds for Figure 17 are selected using the contribution value at the $10^{\text{th}}$ and $50^{\text{th}}$ percentile of these curves.
7.4 Objects and Categories
We offer results for the first two NeRV blocks corresponding to Figure 8 in Figure 19, reserved for this appendix due to space constraints. Overall, the representations for objects are relatively constant over time for Block 2, with some notable exceptions, such as the representation of one of the persons for one of the four sampled neurons, which increases dramatically at the final frame of the video. We also note that Block 1 seems less structured than the other blocks. Specifically, whereas certain object types might dominate for the other blocks, in Block 1 the contributions are more evenly distributed across objects overall.